5. STOICHIOMETRY QUESTION: Silver is widely used for jewelry and for tableware. The tarnish on silver is a layer of silver sulfide that forms according to the following reaction:

4Ag + 2H2S + O2 → 2Ag2S + 2H2O

What mass of silver sulfide would form from the reaction of 0.015 g of silver?
1
Expert's answer
2020-06-17T05:55:57-0400

Solution:

The balanced chemical equation:

4Ag + 2H2S + O2 → 2Ag2S + 2H2

According to the chemical equation: n(Ag)/4 = n(Ag2S)/2

or: n(Ag)/2 = n(Ag2S)


Moles of Ag = n(Ag) = Mass of Ag / Molar mass of Ag

The molar mass of Ag is 107.868 g/mol.

Thus:

n(Ag) = (0.015 g) / (107.868 g/mol) = 0.000139 mol


n(Ag2S) = n(Ag)/2 = (0.000139 mol) / 2 = 0.0000695 mol

Moles of Ag2S = n(Ag2S) = Mass of Ag2S / Molar mass of Ag2S

The Molar mass of Ag2S is 247.8 g/mol.

Thus:

Mass of Ag2S = (0.0000695 mol) × (247.8 g/mol)= 0.017222 g = 0.0172 g


Answer: 0.0172 g of silver sulfide (Ag2S) would form.